# A Bernstein polynomial approach for the estimation of cumulative distribution functions in the presence of missing data

### v1

@@ 4.2 Bandwidth selection via optimized LSCV
where $B(\cdot,\cdot)$ denotes the Beta function (Olver et al. ,, 2010, Section 5.12). Term 1 is evaluated in $\mathcal{O}(m^{2})$ operations.
@@
with $(m_{\min},m_{\mathrm{cap}},c)=(1,300,3)$ by default in our R script. The selected $m^{\star}$ minimizes (4.1).
For the integrated-IPW KDE competitor, the kernel bandwidth $h$ is chosen by an entirely analogous LSCV scheme; see Section 2.2 of Dubnicka, (2009) for details. Obviously, in our setting, we replace the IPW KDE there by its integrated version to get the corresponding CDF estimator (the I-IPW KDE mentioned earlier). Since the optimization and computational structure mirror the Bernstein case, we omit further details.
4.3 Results and discussion
The simulation results are summarized in Table 1 for the pseudo estimators and Table 2 for the feasible estimators. For each sample size $n$, we report the median and interquartile range of the ISEs (top block) and the mean and variance of the ISE (bottom block), each split into Unsmoothed (the IPW empirical CDF), Bernstein (the Bernstein-smoothed IPW empirical CDF), and I-IPW KDE columns.
Pseudo estimators (known propensity)
Across all sample sizes, the integrated-IPW KDE of Dubnicka, (2009) dominates the alternatives. It achieves the smallest mean and median ISE and, simultaneously, the smallest dispersion (interquartile range and variance of the ISE). Bernstein smoothing improves markedly over the unsmoothed IPW empirical CDF (as expected) but remains uniformly behind the I-IPW KDE in this oracle setting.
Feasible estimators (estimated propensity)
When propensities are estimated from the discrete $X$ (the more common case in practice), the ranking depends on the sample size. For small $n$ (e.g., $n=25,50$), the I-IPW KDE yields the lowest mean and median ISE, with interquartile range and variance that are comparable to those of the Bernstein method. For moderate to large $n$ (from about $n\geq 200$; results are mixed for $n=100$), the Bernstein estimator becomes preferable: it attains the smallest mean and median ISE and typically the second smallest dispersion (interquartile range and variance), just behind the unsmoothed IPW empirical CDF. This pattern aligns with our theory for the feasible case, where estimating the propensity scores reduces the leading variance and allows the boundary-adaptive Bernstein smoothing to deliver the best overall risk at larger $n$.
$n$ Median ISE Interquartile Range ISE
Unsmoothed Bernstein I-IPW KDE Unsmoothed Bernstein I-IPW KDE
25 0.0096555 0.0070414 0.0030089 0.0127329 0.0112873 0.0085635
50 0.0047962 0.0043267 0.0011669 0.0075497 0.0084605 0.0037742
100 0.0022661 0.0022514 0.0007979 0.0031720 0.0028093 0.0015168
200 0.0012260 0.0010812 0.0005476 0.0020964 0.0019265 0.0013304
400 0.0006381 0.0005614 0.0002186 0.0009655 0.0008969 0.0005294
800 0.0002860 0.0002664 0.0001286 0.0004353 0.0003667 0.0002541
1600 0.0001638 0.0001569 0.0000879 0.0002642 0.0002915 0.0001743
3200 0.0000647 0.0000625 0.0000458 0.0001016 0.0001147 0.0000600
6400 0.0000350 0.0000326 0.0000207 0.0000523 0.0000504 0.0000335
$n$ Mean ISE Variance ISE ($\times 10^{3}$)
Unsmoothed Bernstein I-IPW KDE Unsmoothed Bernstein I-IPW KDE
25 0.0143201 0.0133532 0.0071434 0.2847269 0.3125405 0.0978971
50 0.0086517 0.0077342 0.0034877 0.0903042 0.0816562 0.0290106
100 0.0031670 0.0029573 0.0015708 0.0101113 0.0098666 0.0043262
200 0.0019346 0.0018317 0.0009948 0.0025334 0.0025783 0.0011975
400 0.0009751 0.0009097 0.0004582 0.0012254 0.0012337 0.0005215
800 0.0004473 0.0004261 0.0002247 0.0002172 0.0002237 0.0000579
1600 0.0002607 0.0002544 0.0001523 0.0000703 0.0000735 0.0000252
3200 0.0001115 0.0001097 0.0000585 0.0000113 0.0000113 0.0000028
6400 0.0000635 0.0000622 0.0000319 0.0000051 0.0000049 0.0000009
Table 1: Pseudo-estimators (known propensity). Top: Median ISE and Interquartile Range ISE. Bottom: Mean ISE and Variance ISE. Columns subdivided into Unsmoothed, Bernstein-smoothed, and Integrated-IPW KDE. Bold indicates the minimum for each $n$ within each group.
$n$ Median ISE Interquartile Range ISE
Unsmoothed Bernstein I-IPW KDE Unsmoothed Bernstein I-IPW KDE
25 0.0036403 0.0032637 0.0019051 0.0043712 0.0069910 0.0057725
50 0.0017426 0.0009735 0.0008677 0.0022932 0.0026344 0.0020441
100 0.0008157 0.0006579 0.0005637 0.0010321 0.0010690 0.0010337
200 0.0004531 0.0003054 0.0004395 0.0005402 0.0005118 0.0007073
400 0.0002851 0.0002310 0.0001647 0.0002725 0.0002801 0.0003617
800 0.0001152 0.0000942 0.0001209 0.0001116 0.0001456 0.0002016
1600 0.0000641 0.0000571 0.0000678 0.0000769 0.0000874 0.0001231
3200 0.0000269 0.0000240 0.0000384 0.0000205 0.0000241 0.0000581
6400 0.0000145 0.0000134 0.0000188 0.0000137 0.0000143 0.0000258
$n$ Mean ISE Variance ISE ($\times 10^{3}$)
Unsmoothed Bernstein I-IPW KDE Unsmoothed Bernstein I-IPW KDE
25 0.0050741 0.0049555 0.0043381 0.0216479 0.0248313 0.0283421
50 0.0025418 0.0019601 0.0016662 0.0046908 0.0042419 0.0038480
100 0.0011946 0.0010309 0.0010398 0.0008924 0.0010181 0.0013722
200 0.0006369 0.0005393 0.0005946 0.0003167 0.0003213 0.0003481
400 0.0003485 0.0002904 0.0003163 0.0000875 0.0000872 0.0001319
800 0.0001436 0.0001235 0.0001630 0.0000121 0.0000104 0.0000191
1600 0.0000848 0.0000793 0.0001094 0.0000038 0.0000040 0.0000099
3200 0.0000348 0.0000316 0.0000493 0.0000007 0.0000007 0.0000019
6400 0.0000173 0.0000165 0.0000268 0.0000001 0.0000001 0.0000005
Table 2: Feasible estimators (estimated propensity). Top: Median ISE and Interquartile Range ISE. Bottom: Mean ISE and Variance ISE. Columns subdivided into Unsmoothed, Bernstein-smoothed, and Integrated-IPW KDE. Bold indicates the minimum for each $n$ within each group.
Overall, the experiments indicate the following practical guidance. If propensities are known, the I-IPW KDE is the strongest performer in our design. In the practically relevant feasible scenario, the I-IPW KDE is attractive at small $n$, while for moderate and large $n$ the Bernstein-smoothed estimator is the recommended choice: it offers the best central accuracy (mean/median ISE) and near-best dispersion.

### v2

@@ 4.2 Smoothing-parameter selection via optimized LSCV
where $B(\cdot,\cdot)$ denotes the Beta function (see, e.g., Olver et al., 2010, Section 5.12). Term 1 is evaluated in $\mathcal{O}(m^{2})$ operations.
@@
with $(m_{\min},m_{\mathrm{cap}},c)=(1,300,5)$ in our R script. The selected $m^{\star}$ minimizes (4.1).
For the I-IPW KDE competitor, the kernel bandwidth $h$ is chosen by an entirely analogous LSCV scheme; see Section 2.2 of Dubnicka (2009) for details.
4.3 Impact of the sample size
This first experiment examines how estimator performance evolves with the sample size $n$ under the benchmark MAR mechanism described above. Tables 1 and 2 report, respectively, the ISE and BISE summaries for the pseudo estimators. Tables 3 and 4 report the analogous quantities for the feasible estimators. Each table reports the mean and standard deviation across replications, with columns ordered as the IPW empirical estimator (labeled Unsmoothed), the I-IPW KDE, and the Bernstein estimator.
$n$ Mean ISE ($\times 10^{8}$) Standard deviation ISE ($\times 10^{8}$)
Unsmoothed I-IPW KDE Bernstein Unsmoothed I-IPW KDE Bernstein
25 1647074 1358550 879985 1608140 1426764 1309825
50 815268 699262 443236 784977 725813 599672
100 391641 340189 224728 427206 399121 341277
200 199960 179626 114586 200241 190795 150616
400 97876 90737 64225 99395 97440 79036
800 48878 46379 38236 52316 52045 42933
1600 24831 23921 24491 25416 25231 22532
3200 12384 12087 12974 13210 13174 12594
6400 6707 6633 6395 6742 6738 6702
Table 1: ISE statistics for pseudo estimators as a function of the sample size.
$n$ Mean BISE ($\times 10^{8}$) Standard deviation BISE ($\times 10^{8}$)
Unsmoothed I-IPW KDE Bernstein Unsmoothed I-IPW KDE Bernstein
25 1139302 1667561 782580 1279327 1685791 1091595
50 517937 894690 406378 596773 876662 537461
100 229159 444997 195663 317125 441765 290511
200 104407 231867 94370 133667 210253 128030
400 54455 125259 51809 72594 116153 71233
800 25206 63876 24492 35856 60658 34790
1600 11926 28213 12104 15886 25092 16289
3200 6052 14266 6047 9102 13351 9078
6400 3049 5101 3078 4236 5406 4278
Table 2: Boundary ISE statistics for pseudo estimators as a function of the sample size.
$n$ Mean ISE ($\times 10^{8}$) Standard deviation ISE ($\times 10^{8}$)
Unsmoothed I-IPW KDE Bernstein Unsmoothed I-IPW KDE Bernstein
25 954524 713278 239168 798381 589142 550044
50 455141 346095 120758 401704 325848 281482
100 220940 174579 67042 200497 172123 143507
200 112125 93299 40544 102896 93909 69979
400 56231 49072 27131 53764 51132 37264
800 26230 23670 18487 21966 21336 14154
1600 13543 12685 13695 11894 11753 8238
3200 6939 6648 7774 6027 5993 5908
6400 3794 3722 3483 3371 3367 3502
Table 3: ISE statistics for feasible estimators as a function of the sample size.
$n$ Mean BISE ($\times 10^{8}$) Standard deviation BISE ($\times 10^{8}$)
Unsmoothed I-IPW KDE Bernstein Unsmoothed I-IPW KDE Bernstein
25 363853 1009706 46069 357238 879450 91124
50 118404 500208 16164 120360 439568 28589
100 38755 265221 6693 36463 194615 3494
200 12347 141440 $33$ 11561 88932 706
400 3997 73016 1682 3858 38645 319
800 1099 37007 797 1080 16500 215
1600 289 17649 311 296 7321 121
3200 57 8144 88 61 2052 48
6400 12 2213 21 12 1271 12
Table 4: Boundary ISE statistics for feasible estimators as a function of the sample size.
Across Tables 14, both the mean and the standard deviation of the error measures decrease steadily with $n$, as expected. For global accuracy, measured by the ISE, the Bernstein estimator clearly dominates in the small-to-medium sample-size range in both the pseudo and feasible regimes: from $n=25$ up to $n=800$, it has the smallest mean ISE and the smallest standard deviation in every reported case. For larger sample sizes, the three procedures become much closer, which is exactly what one would expect since they share the same first-order asymptotics. In particular, in both the pseudo and feasible regimes, the I-IPW KDE has a slightly smaller mean ISE at $n=1600$ and $n=3200$, while the Bernstein estimator is again best at $n=6400$; these differences are numerically very small compared with the pronounced gains seen at smaller $n$. The standard deviations remain smallest for the Bernstein estimator throughout essentially the whole range, up to negligible near-ties at the very largest sample sizes.
For boundary accuracy, the picture is even sharper. The I-IPW KDE is consistently disfavored by the BISE criterion, often by a wide margin, reflecting the familiar boundary spillover of the Gaussian kernel. This boundary effect becomes less damaging as $n$ grows, because the selected bandwidth decreases and the boundary region $0,\delta]\cup[1-\delta,1]$ itself shrinks. Even so, the BISE normalization by $2\delta$ keeps boundary leakage visible, so the KDE remains clearly inferior on that metric. By contrast, the Bernstein estimator is best in the small-to-medium range for both pseudo and feasible estimators, in mean as well as in standard deviation. For large sample sizes, the unsmoothed IPW estimator and the Bernstein estimator become nearly indistinguishable for BISE, with the unsmoothed estimator occasionally having a marginally smaller mean in the pseudo regime ($n=1600$ and $n=6400$) and more systematically in the feasible regime from $n=1600$ onward, whereas Bernstein typically retains the smaller or comparable dispersion.
From a practical standpoint, the small and medium sample sizes are the most relevant, and there the Bernstein estimator wins decisively, both globally and near the boundary.
4.4 Impact of the missing rate
This second experiment examines how estimator performance changes as the missing rate increases. We fix $n=400$ and retain the same latent Gaussian dependence structure between $X$ and $Y$, while varying only the intercept of the logistic MAR mechanism to obtain target missing rates $5\%$, $10\%$, $15\%$, $20\%$, $25\%$, $30\%$, $35\%$, and $40\%$. Tables 5 and 6 report, respectively, the ISE and BISE summaries for the pseudo estimators. Tables 7 and 8 report the analogous quantities for the feasible estimators. Each table reports the mean and standard deviation across replications, with columns ordered as the IPW empirical estimator (labeled Unsmoothed), the I-IPW KDE, and the Bernstein estimator.
Missing rate (%) Mean ISE ($\times 10^{8}$) Standard deviation ISE ($\times 10^{8}$)
Unsmoothed I-IPW KDE Bernstein Unsmoothed I-IPW KDE Bernstein
5 49846 45040 30627 45293 43778 33023
10 59864 54385 37470 54372 52865 40325
15 71677 65463 47369 70809 68961 56305
20 79770 73324 51737 79509 77866 63512
25 101121 93653 67903 105151 102533 87499
30 125709 117153 82657 131487 128308 103824
35 146234 136613 99886 165434 161487 136999
40 176410 163611 121891 193369 188011 159135
Table 5: ISE statistics for pseudo estimators as a function of the missing rate.
Missing rate (%) Mean BISE ($\times 10^{8}$) Standard deviation BISE ($\times 10^{8}$)
Unsmoothed I-IPW KDE Bernstein Unsmoothed I-IPW KDE Bernstein
5 10047 57672 8202 11454 36879 9716
10 16493 65693 14816 19983 50029 18638
15 26482 79045 25054 36023 63341 35503
20 36640 98535 35216 51437 88029 50568
25 52222 116896 49833 69333 109618 68119
30 68140 148497 64454 87625 134854 85735
35 89121 182761 85419 122119 169637 120135
40 115383 209867 112690 154893 206946 154663
Table 6: Boundary ISE statistics for pseudo estimators as a function of the missing rate.
Missing rate (%) Mean ISE ($\times 10^{8}$) Standard deviation ISE ($\times 10^{8}$)
Unsmoothed I-IPW KDE Bernstein Unsmoothed I-IPW KDE Bernstein
5 44112 39257 25455 39272 37832 27858
10 48057 42640 26531 41309 39581 28731
15 51462 45414 27984 46962 44899 33173
20 52249 45862 26533 46526 44178 32603
25 56802 49492 27196 47827 45563 33449
30 62124 53937 28045 56075 52849 38080
35 69724 59972 30044 61193 57088 41029
40 75947 64513 30308 61937 57043 42423
Table 7: ISE statistics for feasible estimators as a function of the missing rate.
Missing rate (%) Mean BISE ($\times 10^{8}$) Standard deviation BISE ($\times 10^{8}$)
Unsmoothed I-IPW KDE Bernstein Unsmoothed I-IPW KDE Bernstein
5 2970 49386 1624 2803 24490 434
10 3126 52148 $16$ 3014 27510 412
15 3387 57372 1629 3273 29030 438
20 3640 66451 1661 3650 36393 357
25 3827 69887 1673 3608 37206 336
30 4187 82539 1700 4416 42584 346
35 4550 94309 1720 4680 48375 381
40 5079 106013 1745 5100 55943 325
Table 8: Boundary ISE statistics for feasible estimators as a function of the missing rate.
Tables 58 show that increasing missingness makes the problem harder, as expected. The mean and standard deviation of the ISE rise with the missing rate for all three methods in both regimes, and the same overall deterioration is visible for BISE as well, especially for the unsmoothed estimator and the I-IPW KDE. The Bernstein estimator is uniformly best across the board: in both the pseudo and feasible regimes, it has the smallest mean and the smallest standard deviation for both ISE and BISE at every reported missing rate. This is fully consistent with the sample-size experiment, since the present design fixes $n=400$, which lies in the range where Bernstein smoothing already showed its clearest advantage.
The boundary-spillover effect of the I-IPW KDE is again very noticeable. For ISE, the KDE typically sits between the unsmoothed and Bernstein estimators, but for BISE it is systematically much worse than the other two procedures, often by a very large margin. The feasible Bernstein estimator is particularly strong: even at $40\%$ missingness, it still has the smallest mean ISE and BISE and the smallest dispersion by a comfortable margin. More generally, while higher missingness degrades performance for all methods, the Bernstein smoother remains the most robust procedure in this experiment, both globally and at the boundary.
4.5 Heuristic order of the LSCV-selected degree
A full proof of the asymptotic behavior of the LSCV selector is beyond the scope of this paper, but the expected scale of $m^{\star}$ is quite clear and is the exact Bernstein analogue of the kernel-CDF argument of Bowman et al. (1998). To keep the notation light, write the oracle version of (4.1) as
$H_{n}(m)\vcentcolon=\int_{0}^{1}\widetilde{F}_{n,m}(y)^{2}\,\mathrm{d}y-\frac{2}{n}\sum_{i=1}^{n}W_{i}\int_{Y_{i}}^{1}\widetilde{F}_{n,m}^{(-i)}(y)\,\mathrm{d}y,\qquad W_{i}=\delta_{i}/\pi(\boldsymbol{X}_{i}),$
and set $R_{n}(m)\vcentcolon=\mathrm{MISE}[\widetilde{F}_{n,m}]$. Since $\mathsf{E}[W_{i}\mathds{1}_{\{Y_{i}\leq y\}}]=F(y)$, since $\widetilde{F}_{n,m}^{(-i)}$ is independent of $(W_{i},Y_{i})$, and since $\mathsf{E}[\widetilde{F}_{n,m}^{(-i)}(y)]=\mathcal{B}_{m}(F)(y)$, we get
$\mathsf{E}[H_{n}(m)]$ $=\int_{0}^{1}\mathsf{E}[\widetilde{F}_{n,m}(y)^{2}]\mathrm{d}y-2\int_{0}^{1}F(y)\mathcal{B}_{m}(F)(y)\mathrm{d}y$
$=\int_{0}^{1}\mathsf{Var}(\widetilde{F}_{n,m}(y))\mathrm{d}y+\int_{0}^{1}\{\mathcal{B}_{m}(F)(y)-F(y)\}^{2}\mathrm{d}y-\int_{0}^{1}F(y)^{2}\mathrm{d}y$
$=R_{n}(m)-\int_{0}^{1}F(y)^{2}\mathrm{d}y.$
Thus, up to the irrelevant constant $\int_{0}^{1}F(y)^{2}\mathrm{d}y$ (which does not depend on $m$), the oracle LSCV curve is an unbiased estimator of the oracle MISE curve.
The key binomial moment identities
$\sum_{k=0}^{m}(k/m-y)\,b_{m,k}(y)=0,\qquad\sum_{k=0}^{m}(k/m-y)^{2}\,b_{m,k}(y)=\frac{y(1-y)}{m}$
show that the binomial basis behaves like a kernel with effective bandwidth of order $m^{-1/2}$. For this reason, one expects the proof of Bowman et al. (1998) to carry over after the substitution $h=m^{-1/2}$. More precisely, one should expand $H_{n}(m)-\mathsf{E}[H_{n}(m)]$ as a linear term, a degenerate quadratic term, and a diagonal term, exactly as in their proof. If
$J_{n}\vcentcolon=\int_{0}^{1}\left(\{\widetilde{F}_{n}(y)-F(y)\}^{2}-\mathsf{E}[\{\widetilde{F}_{n}(y)-F(y)\}^{2}]\right)\mathrm{d}y,$
then the Bernstein analogue of their Theorem 1 should give
$H_{n}(m)+J_{n}=R_{n}(m)+\Delta_{n}(m),$
where, for every $\delta>0$,
$\Delta_{n}(m)=\mathcal{O}\left(n^{\delta}\left\{n^{-1/2}m^{-3/2}+n^{-1}m^{-3/4}+n^{-3/2}\right\}\right)$
with probability tending to 11, uniformly for $m=n^{2/3}u$ with $u$ in any fixed compact subset of $(0,\infty)$. These are exactly the Bernstein counterparts of the terms $n^{-1/2}h^{3}$, $n^{-1}h^{3/2}$, and $n^{-3/2}$ in Equation (4) of Bowman et al. (1998). On the scale $m=n^{2/3}u$, the right-hand side is $o(n^{-4/3})$ as soon as $\delta<1/6$.
Now let
$A_{V}\vcentcolon=\int_{0}^{1}V(y)\mathrm{d}y,\qquad A_{B}\vcentcolon=\int_{0}^{1}B^{2}(y)\mathrm{d}y,$
and note from Corollary 4 that
$R_{n}(m)=n^{-1}\int_{0}^{1}\sigma^{2}(y)\mathrm{d}y-n^{-1}m^{-1/2}A_{V}+m^{-2}A_{B}+\mathrm{o}(n^{-1}m^{-1/2})+\mathrm{o}(m^{-2}).$
If $m=n^{\alpha}$, then
$R_{n}(m)-n^{-1}\int_{0}^{1}\sigma^{2}(y)\mathrm{d}y=A_{B}n^{-2\alpha}-A_{V}n^{-1-\alpha/2}+o\left(n^{-2\alpha}+n^{-1-\alpha/2}\right).$
Hence $\alpha<2/3$ makes the positive bias term dominate, while $\alpha>2/3$ makes the variance gain smaller than the $n^{-4/3}$ improvement available on the balanced scale. Thus the only possible order is $\alpha=2/3$. Writing $m=n^{2/3}u$, we get
$R_{n}(m)=n^{-1}\int_{0}^{1}\sigma^{2}(y)\mathrm{d}y+n^{-4/3}\{\psi(u)+o(1)\},\qquad\psi(u)=A_{B}u^{-2}-A_{V}u^{-1/2},$
and
$\psi^{\prime}(u)=-2A_{B}u^{-3}+\frac{1}{2}A_{V}u^{-3/2}.$
Therefore $\psi$ has the unique minimizer
$u_{0}=\left(\frac{4A_{B}}{A_{V}}\right)^{2/3}=\left[\frac{4\int_{0}^{1}B^{2}(y)\mathrm{d}y}{\int_{0}^{1}V(y)\mathrm{d}y}\right]^{2/3}.$
Fix $\varepsilon>0$. Since $\psi$ has a unique minimum at $u_{0}$, there exists $c_{\varepsilon}>0$ such that $\psi(u)\geq\psi(u_{0})+3c_{\varepsilon}$ whenever $|u-u_{0}|\geq\varepsilon$ and $u$ stays in a fixed compact set containing $u_{0}$. The uniform bound on $\Delta_{n}(m)$ is $o(n^{-4/3})$, so the same inequality holds for $H_{n}(n^{2/3}u)$ with probability tending to 11. This gives the heuristic conclusion
$m^{\star}n^{-2/3}\to u_{0},$
in probability as $n\to\infty$, and in particular
$\mathsf{P}\left(c_{1}n^{2/3}\leq m^{\star}\leq c_{2}n^{2/3}\right)\to 1,$
for any $0<c_{1}<u_{0}<c_{2}<\infty$. For the feasible selector, Corollary 9 shows that propensity estimation only changes the $m$-free $n^{-1}$ term and adds an $m$-dependent remainder $\mathcal{O}(n^{-1}m^{-1})$, which is $o(n^{-4/3})$ when $m$ is of order $n^{2/3}$, so the same asymptotic scale should prevail. Thus the cap $m_{\max}(n)\propto n^{2/3}$ used in Section 4.2 is not ad hoc.
Finally, this heuristic is also in line with what is already known for Bernstein polynomial smoothing in the full-data CDF setting. The expansions in Leblanc (2012a) show that the leading bias and variance terms balance when $m$ is of order $n^{2/3}$, which plays the same heuristic role for Bernstein smoothing that the bandwidth order $n^{-1/3}$ plays for classical kernel CDF smoothing. See also Leblanc (2012b) for boundary-region properties. We are not aware of a published result establishing the analogue of $m^{\star}n^{-2/3}\to u_{0}$ for an LSCV selector in the Bernstein CDF setting. Nevertheless, Dutta (2016) develops data-driven Bernstein estimators with random degree by minimizing estimated pointwise MSE or global MISE and proves consistency and convergence-rate results, with the search for $m$ restricted to intervals proportional to $n^{2/3}$.
Several closely related Bernstein-based results are also worth noting, although none of them proves an asymptotic theorem for an LSCV-selected degree of a Bernstein CDF estimator. In the multivariate bounded-support setting, Wang and Guan (2019) propose a Bernstein polynomial model for multivariate distribution and density estimation, choose the coordinatewise degrees by a change-point rule, and obtain a nearly parametric rate in mean $\chi^{2}$-divergence for the resulting density estimator. For univariate densities on $[0,1]$ that may be unbounded at 0, Bouezmarni and Rolin (2007) establish uniform weak and strong consistency on compact subsets of $(0,1)$ and develop both least-squares and likelihood cross-validation rules for selecting the smoothing parameter. For copula densities that may be unbounded at the corners, Bouezmarni et al. (2013) establish boundary and interior asymptotic properties of the Bernstein estimator and study a least-squares cross-validation rule for the smoothing parameter, with simulations examining finite-sample performance.